# Armijo Line-search Can Make (Stochastic) Gradient Descent Provably Faster

### v2

@@ Abstract
Armijo line-search (Armijo-LS) is a standard method to set the step-size for gradient descent (GD). For smooth functions, Armijo-LS alleviates the need to know the global smoothness constant $L$ and adapts to the “local” smoothness, enabling GD to converge faster. Existing theoretical analyses show that GD with Armijo-LS (GD-LS) can result in constant factor improvements over GD with a $1/L$ step-size (denoted as GD(1/L)). We strengthen these results and show that if the objective function satisfies a certain non-uniform smoothness condition, GD-LS can result in a faster convergence rate than GD(1/L). In particular, we prove that for convex objectives corresponding to logistic regression and multi-class classification, GD-LS can converge to the optimum at a linear rate, and hence improves over the sublinear convergence of GD(1/L). Furthermore, for non-convex objectives satisfying gradient domination (e.g. those corresponding to the softmax policy gradient in RL or generalized linear models with a logistic link function), GD-LS can match the fast convergence of algorithms tailored for these specific settings. Finally, we prove that under the interpolation assumption, for convex losses, stochastic GD with a stochastic line-search can match the fast convergence of GD-LS.
@@ 1 Introduction
Contribution 5. Finally, in Sec. 6, we show that the advantages of line-search carry over to the stochastic setting. Specifically, we consider a finite-sum objective $f(\theta)=\frac{1}{n}\sum_{i=1}^{n}\,f_{i}(\theta)$, and study the convergence of stochastic gradient descent (SGD) in conjunction with a stochastic line-search proposed in Vaswani et al. (2019b). We restrict our attention to the interpolation setting (Vaswani et al., 2019a; Ma et al., 2018; Schmidt & Roux, 2013) which implies that each $f_{i}$ is minimized at $\theta^{*}:=\operatorname*{arg\,min}f(\theta)$. The interpolation assumption is of practical interest (Zhang et al., 2017; Belkin et al., 2019), for example, it is satisfied for logistic regression with linearly separable data.
Interpolation enables the fast convergence of SGD, allowing it to match the GD convergence rate, but with an $O(1)$ iteration cost. Under interpolation, SGD with a stochastic line-search (referred to as SGD-SLS) and its variants empirically outperform constant step-size SGD, and have been used to train deep neural networks (Vaswani et al., 2019b; Galli et al., 2024). We provide further theoretical justification for the fast convergence of SGD-SLS. Specifically, in Cor. 5, we prove that for logistic regression with linearly separable data, SGD-SLS can converge at a linear rate.
@@ 3 GD with Armijo Line-search
The update for gradient descent (GD) with Armijo line-search  (Armijo, 1966) (henceforth referred to as GD-LS) at iteration $t\in[T]$ is given as:
$\theta_{t+1}$ $=\theta_{t}-{\eta_{t}}\,\nabla f(\theta_{t})\,,$ (5)
where ${\eta_{t}}$ is the step-size returned by the backtracking Armijo line-search (referred to as Armijo-LS). In particular, starting from an initial maximum step-size $\eta_{\max}$, Armijo-LS uses backtracking to select the (approximately) largest step-size that satisfies the Armijo condition,
@@
To that end, we first show that, when minimizing non-uniform smooth functions satisfying Assn. 2 and 3, GD-LS can result in provably larger step-sizes that enable faster convergence. For the subsequent theoretical analysis, we only consider “exact backtracking” i.e. we assume that the backtracking procedure returns the largest step-size that satisfies the Armijo condition, meaning that $\beta\approx 1$. Similar to the standard analysis (Nocedal & Wright, 2006), it is straightforward to relax this assumption
We first prove the following lemma that lower-bounds the step-size returned by GD-LS.
@@
If $f$ satisfies Assn. 1, 2 and 3, at iteration $t$, GD-LS returns a step-size
${\eta_{t}}\geq\min\left\{\eta_{\max},\frac{1}{\lambda_{0}+\lambda_{1}\,f(\theta_{t})}\right\}\,,$
where $\lambda_{0}:=3\,\frac{L_{0}+L_{1}\,\omega}{(1-c)}$ and $\lambda_{1}:=3\,\frac{L_{1}(\nu+1)}{(1-c)}$.
Proof Sketch.
For “smaller” step-sizes $\eta\leq\frac{1}{\lambda_{0}+\lambda_{1}\,f(\theta_{t})}$, we use Eq. 1 to prove that the step-size will satisfy the Armijo condition, and the backtracking line-search will terminate. However, since Eq. 1 is only valid when $\left\|y-x\right\|$ is small, we cannot directly use it when the step-size returned by Armijo-LS is “large”. Consequently, we first prove that $g(\theta):=\ln(L_{0}+L_{1}\,f(\theta))$ is $L_{1}$ uniformly-smooth. We then show that the step-size returned by the line-search on $f(\theta)$ satisfies an equivalent Armijo condition on $g(\theta)$. Using the uniform smoothness of $g(\theta)$ along with the standard line-search analysis enables us to lower-bound the step-size. ∎
@@
iterations to ensure to ensure that $f(\theta_{T})-f^{*}\leq\epsilon$.
@@ 4 GD-LS for Convex Losses
For a fixed $\epsilon>0$, assuming $f(\theta)$ is convex and satisfies Assn. 1, 2 and 3 with $L_{0}=0$ and $\omega=0$, GD-LS with $\eta_{\max}=\infty$, requires
$T\geq\max\{2\lambda_{1}\,\left\|\theta_{0}-\theta^{*}\right\|_{2}^{2},1\}\,\left(\frac{f^{*}}{\epsilon}+1\right)\,\ln\left(\frac{f(\theta_{0})-f^{*}}{\epsilon}\right)$
iterations to ensure that $f(\theta_{T})-f^{*}\leq\epsilon$.
@@
While GD(1/L) results in an $\Omega(1/\epsilon)$ rate for general smooth, convex functions (Nesterov et al., 2018), analyses of GD on logistic regression often exploit strong-convexity and prove faster rates (Karimi et al., 2016). In particular, if the iterates lie in a bounded set, the objective is $\mu(\theta)$ strongly-convex where $\mu(\theta)=\lambda_{\min}[X^{T}X]\,\min_{i}\pi_{i}(\theta)\,(1-\pi_{i}(\theta))$ where $\pi_{i}=\frac{1}{1+\exp(-y_{i}\,\langle x_{i},\theta\rangle)}$. Notice that as $\pi_{i}(\theta)$ tends to either zero or one i.e. the predictions become deterministic, $\mu(\theta)$ tends to 00Freund et al. (2018, Theorem 3.3) characterize the resulting rate for GD(1/L) and prove that the suboptimality scales as $O\left(\exp(-T\,\exp\left(\nicefrac{{-1}}{{\xi}}\right))\right)$ where $\xi$ is the degree of non-separability and tends to zero as the data becomes more separable. Hence, the rate becomes exponentially worse as $\xi$ decreases. However, as the data becomes separable, GD-LS converges at a faster linear rate (see Fig. 2), meaning that strong-convexity cannot explain this behaviour.
@@
For an initialization $\theta_{0}$, $\epsilon\in(0,f(\theta_{0}))$ and comparator $u$ s.t. $f(u)\leq\epsilon$, if $f(\theta)$ is convex, satisfies Assn. 1, 2 and 3 with $L_{0}=0,\,\omega=0$, then, GD-LS with $\eta_{\max}=\infty$, $c>\frac{1}{2}$, requires
$T\geq\frac{c\,\lambda_{1}\,\left\|\theta_{0}-u\right\|_{2}^{2}}{\,(2c-1)}\,\left[1+\frac{f(u)}{\epsilon}\right]$
iterations to ensure that $f(\theta_{T})-f(u)\leq\epsilon$.
@@
For logistic regression on linearly separable data with margin $\gamma$, if, for all $i$, $\left\|x_{i}\right\|\leq 1$, for an initialization $\theta_{0}$, an $\epsilon\in(0,f(\theta_{0}))$, GD-LS with $\eta_{\max}=\infty$ requires
$T\geq O\left(\frac{c}{(1-c)\,(2c-1)\,\gamma^{2}}\,\left[\ln\left(\frac{1}{\epsilon}\right)\right]^{2}\right)$
iterations to ensure that $f(\theta_{T})\leq 2\,\epsilon$.
@@
Convergence under $(L_{c},L_{g})$ non-uniform smoothness: Using the reduction in Prop. 3 allows characterizing the convergence of GD-LS on twice-differentiable, non-negative, $(L_{c},L_{g})$ non-uniform smooth (Zhang et al., 2019) and convex functions. In particular, by using Prop. 3 in conjunction with Thm. 1 allows us bound the sub-optimality for the last-iterate of GD-LS (see Cor. 8 in App. C for the formal statement) without requiring the knowledge of $L_{c}$ or $L_{g}$. In particular, assuming $L_{g}\geq 1$, $L_{c}\geq 1$, we prove that GD-LS requires $\tilde{O}\left(\frac{L_{c}\,L_{g}^{2}\,R}{\epsilon}+\frac{L_{g}\,(L_{g}^{2}+L_{c})\,R\,f^{*}}{\epsilon}\right)$ iterations. In contrast, in this same setting, Vankov et al. (2024); Gorbunov et al. (2024) show that GD with the Polyak step-size requires $O\left(\max\left\{\frac{L_{c}}{\epsilon},L_{g}\,R\right\}\right)$ iterations for the best-iterate to achieve an $\epsilon$ sub-optimality (Gorbunov et al., 2024, Thm. 4.1). Comparing the two results, we note that, in general, the upper-bound in (Gorbunov et al., 2024) is tighter in terms of $\epsilon$. However, when $\epsilon=\Theta(f^{*})$, the two bounds are equivalent in their $\epsilon$ dependence. Moreover, while the results in Gorbunov et al. (2024); Vankov et al. (2024) hold for the best-iterate of GD with the Polyak step-size (that requires the knowledge of $f^{*}$), Cor. 8 holds for the last-iterate and does not require knowing $f^{*}$.
@@ 5 GD-LS for Non-convex Losses
$f$ satisfies a non-uniform gradient domination condition if there exists a constant $\zeta\geq 1$ and $\mu(\theta)>0$ s.t. for all $\theta$,
$\left\|\nabla f(\theta)\right\|^{\zeta}\geq\mu(\theta)\,[f(\theta)-f^{*}].$
@@ 5.1 Gradient domination with $\zeta=1$
Given an MAB problem with $K$ arms and known deterministic rewards $r\in[0,1]^{K}$, consider the class of softmax policies $\pi_{\theta}\in\Delta_{K}$ parameterized by $\theta\in\mathbb{R}^{K}$ s.t. $\pi_{\theta}(a)=\frac{\exp(\theta(a))}{\sum_{a^{\prime}}\exp(\theta(a^{\prime}))}$. The loss corresponding to the bandit problem is given by:
$f(\theta)=r(a^{*})-\langle\pi_{\theta},r\rangle\,,$
where $a^{*}:=\operatorname*{arg\,max}_{a\in[K]}r(a)$ is the optimal arm. $f(\theta)$ is non-negative, satisfies Assn. 2 with $L_{0}=0$ and $L_{1}=72$Assn. 3 with $\nu=24$ and $\omega=0$ and Assn. 4 with $\zeta=1$ and $\mu(\theta)=\pi_{\theta}(a^{*})$.
Softmax policy gradient methods (Williams, 1992) optimize the above non-convex objective using gradient descent, and have been analyzed recently (Mei et al., 2020; Agarwal et al., 2021). We aim to use GD-LS to optimize the objective defined in Prop. 4 for which the GD update is given by: $\theta_{t+1}=\theta_{t}-{\eta_{t}}\,\nabla f(\theta_{t})=\theta_{t}+{\eta_{t}}\nabla_{\theta}\langle\pi_{\theta},r\rangle$, and the corresponding Armijo condition is equal to:
$\langle\pi_{\theta_{t+1}},r\rangle$ $\geq\langle\pi_{\theta},r\rangle+c{\eta_{t}}\,\left\|\nabla_{\theta}\langle\pi_{\theta},r\rangle\right\|_{2}^{2}.$
@@
For an $\epsilon>0$, assuming $f(\theta)$ satisfies Assn. 1, 2 and 3 with $L_{0}=0,\,\omega=0$ and Assn. 4 with $\zeta=1$, if $\mu:=\min_{t\in[T]}\mu(\theta_{t})$, then, GD-LS with $\eta_{\max}=\infty$, requires
$T$ $\geq\max\left\{1,\frac{2\,\lambda_{1}}{\mu^{2}}\right\}\,\left(\frac{{f^{*}}}{\epsilon}+1\right)\,\ln\left(\frac{f(\theta_{0})-{f^{*}}}{\epsilon}\right)$
iterations to ensure $f(\theta_{T})\leq\epsilon$.
@@
For an MAB problem with $K$ arms, rewards bounded in $[0,1]$, GD-LS with a uniform initialization i.e. afor-$\forall a$, $\pi_{\theta_{0}}(a)=\nicefrac{{1}}{{K}}$, $c=\frac{1}{2}$, $\eta_{\max}=\infty$ requires
$T=O\left(K^{2}\ln\left(\nicefrac{{1}}{{\epsilon}}\right)\right)$
iterations to guarantee $\langle\pi_{\theta_{T}},r\rangle\geq r(a^{*})-\epsilon$.
Hence, for MAB, GD-LS converges at a linear rate. Under additional assumptions, by combining Prop. 7 and 3, we can prove a similar linear rate for tabular MDPs in the exact setting (see Cor. 9 in Sec. D.1 for the formal result).
The above result is in contrast to GD(1/L) which can only attain an $\Omega\left(\frac{1}{\epsilon}\right)$ convergence rate for both bandits and MDPs (Mei et al., 2020, Theorem 9, 10). The convergence rate of GD-LS matches that of algorithms designed for this specific problem, including GD with a specific line-search that requires the knowledge of $f^{*}$ (Lu et al., 2024), GD with specific increasing step-sizes (Liu et al., 2024), normalized GD (Mei et al., 2021), natural policy gradient (Kakade & Langford, 2002; Xiao, 2022) and mirror descent with a log-sum-exp mirror map (Asad et al., 2024). From a practical perspective, Lu et al. (2024, Figure 1) empirically demonstrate the linear convergence of GD-LS on tabular Markov decision processes, and hence, Cor. 9 substantiates their results theoretically.

### v3

@@ Abstract
Armijo line-search (Armijo-LS) is a standard method to set the step-size for gradient descent (GD). For smooth functions, Armijo-LS alleviates the need to know the global smoothness constant $L$ and adapts to the “local” smoothness, enabling GD to converge faster. Existing theoretical analyses show that GD with Armijo-LS (GD-LS) can result in constant factor improvements over GD with a $1/L$ step-size (denoted as GD(1/L)). We strengthen these results and show that if the objective function satisfies a certain non-uniform smoothness condition, GD-LS can result in a faster convergence rate than GD(1/L). In particular, we prove that for convex objectives corresponding to logistic regression and multi-class classification, GD-LS can converge to the optimum at a linear rate, and hence improves over the sublinear convergence of GD(1/L). Furthermore, for non-convex objectives satisfying gradient domination (e.g. those corresponding to the softmax policy gradient in RL or generalized linear models with a logistic link function), GD-LS can match the fast convergence of algorithms tailored for these specific settings. Finally, we analyze the convergence of stochastic GD with a stochastic line-search on convex losses under the interpolation assumption.
@@ 1 Introduction
Contribution 5. Finally, in Sec. 6, we investigate whether the advantages of line-search carry over to the stochastic setting. Specifically, we consider a finite-sum objective $f(\theta)=\frac{1}{n}\sum_{i=1}^{n}\,f_{i}(\theta)$, and study the convergence of stochastic gradient descent (SGD) in conjunction with a stochastic line-search proposed in Vaswani et al. (2019b). We restrict our attention to the interpolation setting (Vaswani et al., 2019a; Ma et al., 2018; Schmidt & Roux, 2013) which implies that each $f_{i}$ is minimized at $\theta^{*}:=\operatorname*{arg\,min}f(\theta)$. The interpolation assumption is of practical interest (Zhang et al., 2017; Belkin et al., 2019), for example, it is satisfied for logistic regression with linearly separable data.
Interpolation enables the fast convergence of SGD, allowing it to match the GD convergence rate, but with an $O(1)$ iteration cost. Under interpolation, SGD with a stochastic line-search (referred to as SGD-SLS) and its variants empirically outperform constant step-size SGD, and have been used to train deep neural networks (Vaswani et al., 2019b; Galli et al., 2024). We further investigate the convergence of SGD-SLS for logistic regression with linearly separable data, and prove that it can leverage non-uniform smoothness to achieve faster rates.
@@ 3 GD with Armijo Line-search
The update for gradient descent (GD) with Armijo line-search  (Armijo, 1966) (henceforth referred to as GD-LS) at iteration $t\in[T]$ is given as: $\theta_{t+1}=\theta_{t}-{\eta_{t}}\,\nabla f(\theta_{t})$, where $\eta_{t}$} is the step-size returned by the backtracking Armijo line-search (referred to as Armijo-LS). In particular, starting from an initial maximum step-size $\eta_{\max}$, Armijo-LS uses backtracking to select the (approximately) largest step-size that satisfies the Armijo condition,
@@
We first show that, when minimizing non-uniform smooth functions satisfying Assn. 2 and 3, GD-LS can result in provably larger step-sizes that enable faster convergence. For the subsequent theoretical analysis, we only consider “exact backtracking” i.e. we assume that the backtracking procedure returns the largest step-size that satisfies the Armijo condition, meaning that $\beta\approx 1$. Similar to the standard analysis (Nocedal & Wright, 2006), it is straightforward to relax this assumption. We prove the following lemma that lower-bounds the step-size returned by GD-LS.
@@
If $f$ satisfies Assn. 1, 2 and 3, at iteration $t$, GD-LS returns a step-size ${\eta_{t}}\geq\min\left\{\eta_{\max},\frac{1}{\lambda_{0}+\lambda_{1}\,f(\theta_{t})}\right\}$, where $\lambda_{0}:=3\,\frac{L_{0}+L_{1}\,\omega}{(1-c)}$ and $\lambda_{1}:=3\,\frac{L_{1}(\nu+1)}{(1-c)}$.
@@
iterations to ensure that $f(\theta_{T})-f^{*}\leq\epsilon$.
@@ 4 GD-LS for Convex Losses
For a fixed $\epsilon>0$, assuming $f(\theta)$ is convex and satisfies Assn. 1, 2 and 3 with $L_{0}=0$ and $\omega=0$, GD-LS with $\eta_{\max}=\infty$, requires $T\geq\max\{2\lambda_{1}\,\left\|\theta_{0}-\theta^{*}\right\|_{2}^{2},1\}\,\left(\frac{f^{*}}{\epsilon}+1\right)\,\ln\left(\frac{f(\theta_{0})-f^{*}}{\epsilon}\right)$ iterations to ensure that $f(\theta_{T})-f^{*}\leq\epsilon$.
@@
While GD(1/L) results in an $\Omega(1/\epsilon)$ rate for general smooth, convex functions (Nesterov et al., 2018), analyses of GD on logistic regression often exploit strong-convexity and prove faster rates (Karimi et al., 2016). In particular, if the iterates lie in a bounded set, the objective is $\mu(\theta)$ strongly-convex where $\mu(\theta)=\lambda_{\min}[X^{T}X]\,\min_{i}\pi_{i}(\theta)\,(1-\pi_{i}(\theta))$ where $\pi_{i}=\frac{1}{1+\exp(-y_{i}\,\langle x_{i},\theta\rangle)}$. Notice that as $\pi_{i}(\theta)$ tends to either zero or one i.e. the predictions become deterministic, $\mu(\theta)$ tends to 0Freund et al. (2018, Theorem 3.3) characterize the resulting rate for GD(1/L) and prove that the suboptimality scales as $O\left(\exp(-T\,\exp\left(\nicefrac{{-1}}{{\xi}}\right))\right)$ where $\xi$ is the degree of non-separability and tends to zero as the data becomes more separable. Hence, the rate becomes exponentially worse as $\xi$ decreases. However, as the data becomes separable, GD-LS converges at a faster linear rate (see Fig. 2), meaning that strong-convexity cannot explain this behaviour.
@@
For an initialization $\theta_{0}$, $\epsilon\in(0,f(\theta_{0}))$ and comparator $u$ s.t. $f(u)\leq\epsilon$, if $f(\theta)$ is convex, satisfies Assn. 1, 2 and 3 with $L_{0}=0,\,\omega=0$, then, GD-LS with $\eta_{\max}=\infty$, $c>\frac{1}{2}$, requires $T\geq\frac{c\,\lambda_{1}\,\left\|\theta_{0}-u\right\|_{2}^{2}}{\,(2c-1)}\,\left[1+\frac{f(u)}{\epsilon}\right]$ iterations to ensure that $f(\theta_{T})-f(u)\leq\epsilon$.
@@
For logistic regression on linearly separable data with margin $\gamma$, if, for all $i$, $\left\|x_{i}\right\|\leq 1$, for an initialization $\theta_{0}$, an $\epsilon\in(0,f(\theta_{0}))$, GD-LS with $\eta_{\max}=\infty$ requires $T\geq O\left(\frac{c}{(1-c)\,(2c-1)\,\gamma^{2}}\,\left[\ln\left(\frac{1}{\epsilon}\right)\right]^{2}\right)$ iterations to ensure that $f(\theta_{T})\leq 2\,\epsilon$.
@@
Convergence under $(L_{c},L_{g})$ non-uniform smoothness: Using the reduction in Prop. 3 allows characterizing the convergence of GD-LS on twice-differentiable, non-negative, $(L_{c},L_{g})$ non-uniform smooth (Zhang et al., 2019) and convex functions. In particular, by using Prop. 3 in conjunction with Thm. 1 allows us bound the sub-optimality for the last-iterate of GD-LS (see Cor. 5 in App. C for the formal statement) without requiring the knowledge of $L_{c}$ or $L_{g}$. In particular, assuming $L_{g}\geq 1$, $L_{c}\geq 1$, we prove that GD-LS requires $\tilde{O}\left(\frac{L_{c}\,L_{g}^{2}\,R}{\epsilon}+\frac{L_{g}\,(L_{g}^{2}+L_{c})\,R\,f^{*}}{\epsilon}\right)$ iterations. In contrast, in this same setting, Vankov et al. (2024); Gorbunov et al. (2024) show that GD with the Polyak step-size requires $O\left(\max\left\{\frac{L_{c}}{\epsilon},L_{g}\,R\right\}\right)$ iterations for the best-iterate to achieve an $\epsilon$ sub-optimality (Gorbunov et al., 2024, Thm. 4.1). Comparing the two results, we note that, in general, the upper-bound in (Gorbunov et al., 2024) is tighter in terms of $\epsilon$. However, when $\epsilon=\Theta(f^{*})$, the two bounds are equivalent in their $\epsilon$ dependence. Moreover, while the results in Gorbunov et al. (2024); Vankov et al. (2024) hold for the best-iterate of GD with the Polyak step-size (that requires the knowledge of $f^{*}$), Cor. 5 holds for the last-iterate and does not require knowing $f^{*}$.
@@ 5 GD-LS for Non-convex Losses
$f$ satisfies a non-uniform gradient domination condition if there exists a constant $\zeta\geq 1$ and $\mu(\theta)>0$ s.t. for all $\theta$, $\left\|\nabla f(\theta)\right\|^{\zeta}\geq\mu(\theta)\,[f(\theta)-f^{*}]$.
@@ 5.1 Gradient domination with $\zeta=1$
Given an MAB problem with $K$ arms and known deterministic rewards $r\in[0,1]^{K}$, consider the class of softmax policies $\pi_{\theta}\in\Delta_{K}$ parameterized by $\theta\in\mathbb{R}^{K}$ s.t. $\pi_{\theta}(a)=\frac{\exp(\theta(a))}{\sum_{a^{\prime}}\exp(\theta(a^{\prime}))}$. The loss corresponding to the bandit problem is given by: $f(\theta)=r(a^{*})-\langle\pi_{\theta},r\rangle$, where $a^{*}:=\operatorname*{arg\,max}_{a\in[K]}r(a)$ is the optimal arm. $f(\theta)$ is non-negative, satisfies Assn. 2 with $L_{0}=0$ and $L_{1}=72$Assn. 3 with $\nu=24$ and $\omega=0$ and Assn. 4 with $\zeta=1$ and $\mu(\theta)=\pi_{\theta}(a^{*})$.
Softmax policy gradient methods (Williams, 1992) optimize the above non-convex objective using gradient descent, and have been analyzed recently (Mei et al., 2020; Agarwal et al., 2021). We aim to use GD-LS to optimize the objective defined in Prop. 4 for which the GD update is given by: $\theta_{t+1}=\theta_{t}-{\eta_{t}}\,\nabla f(\theta_{t})=\theta_{t}+{\eta_{t}}\nabla_{\theta}\langle\pi_{\theta},r\rangle$, and the corresponding Armijo condition is given as $\langle\pi_{\theta_{t+1}},r\rangle\geq\langle\pi_{\theta},r\rangle+c{\eta_{t}}\,\left\|\nabla_{\theta}\langle\pi_{\theta},r\rangle\right\|_{2}^{2}$.
@@
For an $\epsilon>0$, assuming $f(\theta)$ satisfies Assn. 1, 2 and 3 with $L_{0}=0,\,\omega=0$ and Assn. 4 with $\zeta=1$, if $\mu:=\min_{t\in[T]}\mu(\theta_{t})$, then, GD-LS with $\eta_{\max}=\infty$, requires $T\geq\max\left\{1,\frac{2\,\lambda_{1}}{\mu^{2}}\right\}\,\left(\frac{{f^{*}}}{\epsilon}+1\right)\,\ln\left(\frac{f(\theta_{0})-{f^{*}}}{\epsilon}\right)$ iterations to ensure $f(\theta_{T})\leq\epsilon$.
@@
For an MAB problem with $K$ arms, rewards bounded in $[0,1]$, GD-LS with a uniform initialization i.e. $\forall a$, $\pi_{\theta_{0}}(a)=\nicefrac{{1}}{{K}}$, $c=\frac{1}{2}$, $\eta_{\max}=\infty$ requires $T=O\left(K^{2}\ln\left(\nicefrac{{1}}{{\epsilon}}\right)\right)$ iterations to guarantee $\langle\pi_{\theta_{T}},r\rangle\geq r(a^{*})-\epsilon$.
Hence, for MAB, GD-LS converges at a linear rate. Under additional assumptions, by combining Prop. 7 and 3, we can prove a similar linear rate for tabular MDPs in the exact setting (see Cor. 6 in Sec. D.1 for the formal result).
The above result is in contrast to GD(1/L) which can only attain an $\Omega\left(\frac{1}{\epsilon}\right)$ convergence rate for both bandits and MDPs (Mei et al., 2020, Theorem 9, 10). The convergence rate of GD-LS matches that of algorithms designed for this specific problem, including GD with a specific line-search that requires the knowledge of $f^{*}$ (Lu et al., 2024), GD with specific increasing step-sizes (Liu et al., 2024), normalized GD (Mei et al., 2021), natural policy gradient (Kakade & Langford, 2002; Xiao, 2022) and mirror descent with a log-sum-exp mirror map (Asad et al., 2024). From a practical perspective, Lu et al. (2024, Figure 1) empirically demonstrate the linear convergence of GD-LS on tabular Markov decision processes, and hence, Cor. 6 substantiates their results theoretically.